OpenStudy (anonymous):

How do I find the derivative of f(x)=-10x with the limit process? It's a constant slope...so I don't understand how I would find a secant and tangent line....

OpenStudy (amistre64):

found it

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ F(x+h) - F(x)}{ h }\]

OpenStudy (amistre64):

\[\lim_{h\to 0}\frac{f(x+h)-f(x)}{(x+h)-x}\] \[\lim_{h\to 0}\frac{-10(x+h)-(-10x)}{h}\] \[\lim_{h\to 0}\frac{\cancel{-10x}-10h+\cancel{10x}}{h}\] \[\lim_{h\to 0}\frac{-10\cancel h}{\cancel h}\] \[\lim_{h\to 0}~-10\]

OpenStudy (anonymous):

With a quadratic function, I was able to do this: in order to find the secant line.Now i cant to that because it's linear

OpenStudy (amistre64):

lol, the slope of the line at any given point is ... well, the slope of the line between any 2 point

OpenStudy (amistre64):